OpenStudy (anonymous):

Solve 7 In(3x+2)-8=6

OpenStudy (unklerhaukus):

add 8 to both sides

OpenStudy (anonymous):

7 In(3x+2)=14

OpenStudy (unklerhaukus):

now divide both sides by 7

OpenStudy (anonymous):

in(3x+2)=2

OpenStudy (anonymous):

@enchanted_bubbles

OpenStudy (alphabeta):

Guess you could calculate it like this? (Anyone, please correct me if I'm wrong.) 7ln(3x+2)-8=6 7ln(3x+2)=14 ln(3x+2)=2 e^ln(3x+2)=e^2 3x+2=e^2 3x=(e^2)-2 x= ((e^2)-2)/3 x=1.79635... (value I got on my calculator)

OpenStudy (anonymous):

Thanks :)

OpenStudy (unklerhaukus):

to check our answer:\[7\ln(3x+2)-8=6\\ 7\ln\left(3(\tfrac{e^2-2}3)+2\right)-8=6\\ 7\ln\left(e^2-2+2\right)-8=6\\ 7\ln\left(e^2\right)-8=6\\ 7(2)\ln\left(e\right)-8=6\\ 14-8=6\\ 6=6\\ \checkmark\]